OpenStudy (anonymous):

It is given that x varies directly as y and inversely as the square of z, and that x = 10 when y = 4 and z = 14. What is x when y = 14 and z = 7?

OpenStudy (anonymous):

Typo on last sentence. Supposed to be: what is x when y = 16 and z = 7

OpenStudy (goformit100):

I @goformit100 an OS User, Welcomes you to Open Study. 'Discover. Learn. Earn a Certificate Ask questions, give help, and get certified with 150,000 students from 180 countries. ALL THE BEST.

OpenStudy (anonymous):

x = k (y/z^2) where k is a constant then find the k by substituting x = 10, y = 4 and z = 7 after you got the k value, x = k (16/49)

terenzreignz (terenzreignz):

Variation... roughly speaking, direct variation involves division while inverse variation involves multiplication... So since x varies directly with y, divide x by y, and since x varies inversely with the *square* of z, then multiply x by \(z^2\) And all that should remain constant: \[\Large \frac{\color{red}x\color{blue}z^2}{\color{green}y}= k\]

terenzreignz (terenzreignz):

Now, you need to find the value of k. Use the second bit of your problem: "...and that x = 10 when y = 4 and z = 14"